# TopoLM: brain-like spatio-functional organization in a topographic language model

### v1

@@ Abstract
Neurons in the brain are spatially organized such that neighbors on tissue often exhibit similar response profiles. In the human language system, experimental studies have observed clusters for syntactic and semantic categories, but the mechanisms underlying this functional organization remain unclear. Here, building on work from the vision literature, we develop TopoLM, a transformer language model with an explicit two-dimensional spatial representation of model units. By combining a next-token prediction objective with a spatial smoothness loss, representations in this model assemble into clusters that correspond to semantically interpretable groupings of text and closely match the functional organization in the brain’s language system. TopoLM successfully predicts the emergence of the spatio-functional organization of a cortical language system as well as the organization of functional clusters selective for fine-grained linguistic features empirically observed in human cortex. Our results suggest that the functional organization of the human language system is driven by a unified spatial objective, and provide a functionally and spatially aligned model of language processing in the brain.
@@ 1 Introduction
Artificial neural network (ANN) models of language have recently been shown to accurately predict neural activity in the human language system (Schrimpf et al., 2021b; Caucheteux & King, 2022; Goldstein et al., 2022). When presented with the same text input, the unit activity at internal layers of especially transformer-based models (Vaswani et al., 2017; Radford et al., 2019) is strikingly similar to the internal activity measured experimentally in human cortex. The most powerful models predict even close to 100% of the explainable variance of neural responses to sentences in some brain datasets (Schrimpf et al., 2021a). However, while there is a strong alignment to the brain’s functional responses, a crucial element of cortex is entirely lacking from today’s language models: the spatial arrangement of neurons on the cortical surface.
@@
The topography of the human language system on the other hand lacks a comprehensive computational explanation. Neuroscience experiments suggest both a macro-organization at the level of a distributed cortical network that selectively responds to linguistic processing (Fedorenko et al., 2010; 2011; 2024b; Blank et al., 2014), as well as a micro-organization into clusters that correspond to syntactic and semantic categories such as verbs, nouns, and concrete words (Shapiro et al., 2006; Moseley & Pulvermüller, 2014; Hauptman et al., 2024). What are the mechanisms underlying this spatio-functional organization of the language system in the brain?
Here, we develop TopoLM, a neural network model for brain-like topographic language processing. TopoLM is based on the transformer architecture but incorporates an explicit spatial arrangement of units. We train the model via a combined task and spatial loss, which optimizes the model to perform autoregressive language modeling while encouraging local correlation, similarly to a recent approach used in vision (Lee et al., 2020; Margalit et al., 2024). The spatio-functional organization that emerges in this model is semantically interpretable and aligned with clusters that have been observed experimentally in brain recordings. Comparing TopoLM with a non-topographic baseline model (i.e. one trained without spatial loss) on a series benchmarks, we show that while TopoLM achieves slightly lower scores on some behavioral tasks (BLiMP), its performance on other downstream tasks (GLUE) and on brain alignment benchmarks (using the Brain-Score platform) is on par with the non-topographic control.
@@
Figure 1: Building a topographic language model with a brain-like spatio-functional organization. (a) TopoLM modifies the Transformer architecture with a two-dimensional spatial encoding at the output of each attention and MLP layer. This representation enables the propagation of a spatial correlation loss that encourages smooth response profiles in adjacent units. This spatial loss is jointly optimized with the task loss (autoregressive cross-entropy) during training. (b) At each forward pass, we randomly select five neighborhoods in each layer (of which we here only show 3 for clearity) and compute the pairwise correlation of unit activations within each neighborhood. The spatial loss is computed by comparing these correlations to the inverse distances between associated unit pairs, with the final loss averaged across units pairs and neighborhoods. Computing the loss on cortical neighborhoods is an efficient approximation of the spatial loss. (c) We use a FWHM filter to simulate the fMRI sampling process such that a simulated voxel’s response (‘fMRI-like signal’) is composed of a combination of responses from neighboring units (Kriegeskorte et al., 2010). We simulate the response as a Gaussian random variable, with FWHM $2.0$ mm, assuming unit distances of $1.0$ mm.
@@ 2 Related Work
This spatial organization of the primate visual cortex has prompted work on topographic artificial neural networks for vision. First approaches focused on the organization of inferotemporal cortex, restricting topographic organization to later model layers (TDANN, Lee et al. 2020; ITN, Blauch et al. 2022; DNN-SOM, Zhang et al. 2021, Doshi & Konkle 2023). Recent models are designed such that all layers are topographic and thus mimic the topographic features across the visual cortex—for example, smoothly varying orientation preference maps forming pinwheels in model V1 and category-selective regions in model IT (All-TNNs, Lu et al. 2023; new version of TDANN, Margalit et al. 2024).
@@
Comparatively little work has explored the idea of inducing topography in language models. In particular, the only topographic language model we are aware of is BinHuraib et al. (2024)’s Topoformer, which induces spatial organization onto a single-headed attention Transformer architecture using local connectivity constraints. This model arranges keys and queries on 2D grids, combined with a locally connected layer in the attention mechanism as opposed to full connectivity. Our approach primarily differs from Topoformer in that we use a spatial smoothness loss term to drive the emergence of local correlations, similarly to Lee et al. (2020) and Margalit et al. (2024)’s TDANN vision models. In this sense, our model extends Margalit et al. (2024)’s unifying principle of functional organization from the visual cortex into the language system. TopoLM is thus able to benefit from full connectivity, rather than requiring local connectivity to develop clustering. Because we apply this loss to the output of entire attention mechanism at each layer (as well as to the MLP), TopoLM can also benefit from multi-head attention, which was not explored in Topoformer. Finally, our model also uses an autoregressive task loss, rather than a masked autoencoder objective (as used in BinHuraib et al. (2024)), which has been shown to have higher performance on neural alignment benchmarks (e.g. Schrimpf et al., 2021b).
@@ 3 Model Design and Visualization
Contrary to the convolutional neural network architecture used in topographic vision models (Margalit et al., 2024), we use the Transformer architecture (Vaswani et al., 2017) which is dominant in language modeling. We augment the objective function with a spatial correlation loss, in addition to the cross-entropy task loss. This loss function measures spatial smoothness, which serves as an efficiently computable proxy for neural wiring length: neurons located close to one another should have similar response profiles—i.e. their activations should be correlated (Lee et al., 2020).
To introduce a notion of ‘space’ in the model, we encode the units of each attention and MLP layer in a square grid with a one-to-one mapping. We randomly permute these positions for each layer such that each layer has a unique spatial encoding. On each forward pass, we first compute the pairwise Pearson’s correlation vector $\mathbf{r}_{k}$ between unit activations on the input batch for each layer $k$ (see Figure 1B). If a layer has $N$ units, $\mathbf{r}$ is of dimension $\mathrm{nCr}(N,2)$. Then, the spatial loss for layer $k$ is given by
@@
where $\mathbf{d}_{k}$ is a vector of pairwise inverse distances between units, based on their spatial encoding, and $\mathrm{corr}$ is Pearson’s $r$. This means that nearby units (those with high pairwise inverse distance) should have highly correlated activations on the same inputs, and that distant units should have low correlation; this gives us a notion of spatial smoothness. Note that we scale by a factor of $0.5$ to ensure that $\mathrm{SL}\in[0,1]$.
We compute this spatial loss for every attention and MLP layer in a Transformer, prior to normalization and addition into the residual stream. For each batch of inputs, the model is then optimized subject to the loss criterion
@@
In the below experiments, we utilize an adapted /-2-small style architecture (Radford et al., 2019). We use hidden dimension 784 such that we can evenly embed units in a $28\times 28$ grid. The model has 12 Transformer blocks, each with 16 attention heads and a GELU activation function. We train our models on a randomly sampled 10B-token subset of the FineWeb-Edu dataset. The task loss is cross-entropy on next-word prediction. We use batch size 48 and block size 1024. For spatial loss, we set $\alpha_{k}=2.5$ across all layers [^1] For spatial loss, we set $\alpha_{k}=2.5$ across all layers. We chose this value of $\alpha$ after extensive hyperparameter search. In particular, lower values of $\alpha$ do not adequately encourage the development of topography, while greater values impede task performance and the development of meaningful representations. and operationalize the inverse distance vector $\mathbf{d}$ with the $\ell^{\infty}$ norm. Rather than computing the spatial loss for the entire layer, which is computationally expensive, as in Margalit et al. (2024) we approximate the loss at every batch using small neighborhoods. For each batch, we average the spatial loss across $5$ randomly selected neighborhoods, each of $\ell^{\infty}$ radius $5$. This allows us to compute loss more efficiently without significant performance drops.
@@
Figure 2: Brain-like response profiles across the core language system. (a) Applying a functional localizer (Fedorenko et al., 2010) we isolate the core language system of TopoLM, and find clear brain-like spatial organization (for brevity, we only show Transformer blocks 11-14 here). Response profile across individual language-selective clusters in TopoLM are similar to one another, consistent with (b) the language system in human cortex (Fedorenko et al., 2024b). (c) Across the entire core language system, TopoLM (blue) mostly matches the neural data (green), but not exactly; however, the non-topographic baseline model (orange) fails to capture neural patterns as well.
@@ 4 Spatio-Functional Organization of the Core Language System
Language processing in the brain engages a set of left-lateralized frontal and temporal brain regions. These areas are typically referred to as the ’core language system’ (Fedorenko et al., 2010) and respond selectively to linguistic input in contrast to non-linguistic stimuli (see Fedorenko et al., 2024a, for an overview). Due to anatomical differences between individuals, the language system is defined via a functional localizer that contrasts syntactically and semantically valid sentences against a perceptually matched control, such as strings of nonwords (Fedorenko et al., 2011).
@@
Prior work on neural alignment in language models typically compares neural responses across the entire core language system of the brain to model activations. The topographic organization of our model enables us to test for the emergence of a brain-like spatially organized system core language system in silico. A successful spatio-functional alignment between brain and model would mean that (1) distinct language-selective clusters emerge in the model, (2) these clusters all have consistent response profiles, and (3) the response profiles match the activity profiles in the brain (sentences >>> {unconnected words, jabberwocky} >>> nonwords; AlKhamissi et al., 2024).
@@ Methods.
To isolate the core language system in TopoLM, we use the same localization stimuli as AlKhamissi et al. (2024); Fedorenko et al. (2010), which consists of a set of 160 sentences and 160 strings of non-words, all 12 words each. After passing these through the model, at each attention and MLP layer we run a $t$-test across the activations of all layer units. We then define the core language system as all units that are significantly language-selective ($p<0.05$ after correction for multiple comparison across all layers using the false-discovery-rate (Benjamini & Hochberg, 1995)).
We then define language-selective clusters using an evolutionary clustering algorithm applied to each contrast map. In each layer, we begin with the most selective unit by $t$-value, and then repeatedly add the most selective neighboring unit to the cluster, until we hit a pre-determined $p$-value threshold associated with the unit’s selectivity (here, $p>0.05$). We repeat this process, searching for new clusters until all units in the layer are exhausted. We discard clusters with fewer than $10$ units.
@@ Results.
TopoLM exhibits clear brain-like spatial organization of the language network, such that (1) multiple language-selective clusters emerge across the topographic tissue (Figure 2A and Figure 7), (2) across most clusters, the response profiles are consistent with one another (Figure 8), and (3) response profiles mostly match the ones in the brain. The response profiles are not a perfect match to the brain data—while sentences have higher activations than Jabberwocky and nonword stimuli, they do not have higher activation than unconnected words as in brain data. However, looking across the entire language selective network, the response profile of the non-topographic baseline model similarly fails to capture the neural response profile (Figure 2C), suggesting a general shortcoming of the base transformer model, rather than a weakness of topography.
5 Spatio-Functional Organization Of Syntactic and Semantic Clusters
Beyond selectivity for language in general, experimental evidence supports the existence of cortical noun- and verb-selective clusters across human subjects during processing of verbal and nominal stimuli in auditory (Elli et al., 2019; Hauptman et al., 2024), visual (Moseley & Pulvermüller, 2014), and production (Shapiro et al., 2006) tasks. Here, we compare spatial activation patterns predicted by TopoLM to two groups of fMRI studies: Elli et al. (2019) / Hauptman et al. (2024), who use the same set of stimuli and the same experimental setup (Figure 3 and Figures 6 and 9), and Moseley & Pulvermüller (2014) (Figure 4). We evaluate TopoLM predictions quantitatively using the fMRI data from (Hauptman et al., 2024) and perform qualitative evaluations where no neural data was available (Elli et al., 2019; Moseley & Pulvermüller, 2014).

### v2

@@ Abstract
Neurons in the brain are spatially organized such that neighbors on tissue often exhibit similar response profiles. In the human language system, experimental studies have observed clusters for syntactic and semantic categories, but the mechanisms underlying this functional organization remain unclear. Here, building on work from the vision literature, we develop TopoLM, a transformer language model with an explicit two-dimensional spatial representation of model units. By combining a next-token prediction objective with a spatial smoothness loss, representations in this model assemble into clusters that correspond to semantically interpretable groupings of text and closely match the functional organization in the brain’s language system. TopoLM successfully predicts the emergence of a spatially organized cortical language system as well as the organization of functional clusters selective for fine-grained linguistic features empirically observed in human cortex. Our results suggest that the functional organization of the human language system is driven by a unified spatial objective, and provide a functionally and spatially aligned model of language processing in the brain. [^1] Code available at https://github.com/epflneuroailab/topolm.
@@ 1 Introduction
Artificial neural network (ANN) models of language have recently been shown to accurately predict neural activity in the human language system (Schrimpf et al., 2021; Caucheteux & King, 2022; Goldstein et al., 2022). When presented with the same text input, the unit activity at internal layers of especially transformer-based models (Vaswani et al., 2017; Radford et al., 2019) is strikingly similar to the internal activity measured experimentally in human cortex. The most powerful models predict even close to 100% of the explainable variance of neural responses to sentences in some brain datasets (Schrimpf et al., 2021). However, while there is a strong alignment to the brain’s functional responses, a crucial element of cortex is entirely lacking from today’s language models: the spatial arrangement of neurons on the cortical surface.
@@
The topography of the human language system on the other hand lacks a comprehensive computational explanation. Neuroscience experiments suggest both a macro-organization at the level of a distributed cortical network that selectively responds to linguistic processing (Fedorenko et al., 2010; 2011; 2024; Blank et al., 2014), as well as a micro-organization into clusters that correspond to syntactic and semantic categories such as verbs, nouns, and concrete words (Shapiro et al., 2006; Moseley & Pulvermüller, 2014; Hauptman et al., 2024). What are the mechanisms underlying this spatio-functional organization of the language system in the brain?
Here, we develop TopoLM, a neural network model for brain-like topographic language processing. TopoLM is based on the transformer architecture but incorporates an explicit spatial arrangement of units. We train the model via a combined task and spatial loss, which optimizes the model to perform autoregressive language modeling while encouraging local correlation, similarly to a recent approach used in vision (Lee et al., 2020; Margalit et al., 2024). The spatio-functional organization that emerges in this model is semantically interpretable and aligned with clusters that have been observed experimentally in brain recordings. Comparing TopoLM with a non-topographic baseline model (i.e. one trained without spatial loss) on a series of benchmarks, we show that while TopoLM achieves slightly lower scores on some behavioral tasks (BLiMP), its performance on other downstream tasks (GLUE) and on brain alignment benchmarks (using the Brain-Score platform) is on par with the non-topographic control.
@@
Figure 1: Building a topographic language model with brain-like spatio-functional organization. (a) TopoLM modifies the Transformer architecture with a two-dimensional spatial encoding at the output of each attention and MLP layer. This representation enables the use of a spatial correlation loss that encourages smooth response profiles in adjacent units. This spatial loss is jointly optimized with cross-entropy task loss during training. (b) At each forward pass, we randomly select five neighborhoods in each layer (of which we here only show 3 for clearity) and compute the pairwise correlation of unit activations within each neighborhood. The spatial loss is computed by comparing these correlations to the inverse distances between associated unit pairs, with the final loss averaged across units pairs and neighborhoods. Computing the loss on cortical neighborhoods is an efficient approximation of the spatial loss. (c) We use a FWHM filter to simulate the fMRI sampling process such that a simulated voxel’s response (‘fMRI-like signal’) is composed of a combination of responses from neighboring units (Kriegeskorte et al., 2010). We simulate the response as a Gaussian random variable, with FWHM $2.0$ mm, assuming unit distances of $1.0$ mm.
@@ 2 Related Work
This spatial organization of the primate visual cortex has prompted work on topographic ANNs for vision. First approaches focused on the organization of inferotemporal cortex, restricting topographic organization to later model layers (TDANN, Lee et al. 2020; ITN, Blauch et al. 2022; DNN-SOM, Zhang et al. 2021, Doshi & Konkle 2023). Recent models are designed such that all layers are topographic and thus mimic topographic features across the visual cortex—for example, smoothly varying orientation preference maps forming pinwheels in model V1 and category-selective regions in model IT (All-TNNs, Lu et al. 2023; new version of TDANN, Margalit et al. 2024).
@@
Comparatively little work has explored the idea of inducing topography in language models. In particular, the only topographic language model we are aware of is BinHuraib et al. (2024)’s Topoformer, which induces spatial organization onto a single-headed attention Transformer architecture using local connectivity constraints. This model arranges keys and queries on 2D grids, combined with a locally connected layer in the attention mechanism as opposed to full connectivity.
Our approach primarily differs from Topoformer in that we use a spatial smoothness loss term to drive the emergence of local correlations, similarly to Lee et al. (2020) and Margalit et al. (2024)’s TDANN vision models. In this sense, our model extends Margalit et al. (2024)’s unifying principle of functional organization from the visual cortex into the language system. TopoLM is thus able to benefit from full connectivity, rather than requiring local connectivity to develop clustering. Because we apply this loss to the output of entire attention mechanism at each layer (as well as to the MLP), TopoLM can also benefit from multi-head attention, which empirically improves fits to neural data (AlKhamissi et al., 2024); this was not explored in (BinHuraib et al., 2024). Finally, our model also uses an autoregressive task loss, rather than a masked autoencoder objective (as used in BinHuraib et al. (2024)), which has been shown to have higher performance on neural alignment benchmarks (e.g. Schrimpf et al., 2021).
@@ 3 Model Design and Visualization
Instead of the convolutional neural network architecture used in topographic vision models (Margalit et al., 2024), we use the Transformer architecture (Vaswani et al., 2017) which is dominant in language modeling. We augment the objective function with a spatial correlation loss, in addition to the cross-entropy task loss. This loss function measures spatial smoothness, which serves as an efficiently computable proxy for neural wiring length: neurons located close to one another should have similar response profiles—i.e. their activations should be correlated (Lee et al., 2020).
To introduce a notion of ‘space’ in the model, we bijectively map the units of each attention layer and MLP to a square grid. We randomly permute these positions for each layer such that each layer has a unique spatial encoding. [^2] Our goal is to abstract away from feed-forward propagation as much as possible, as the hierarchical organization of the brain is quite different from that of a language model. This random permutation prevents the model from exploiting the feed-forward nature of the Transformer. Without it, the model minimizes spatial loss by propagating the same spatial pattern through the network; see Figure 12. On each forward pass, we first compute the pairwise Pearson’s correlation vector $\mathbf{r}_{k}$ between unit activations on the input batch for each layer $k$ (see Figure 1B). If a layer has $N$ units, $\mathbf{r}$ is of dimension $\mathrm{nCr}(N,2)$. Then, the spatial loss for layer $k$ is given by
@@
where $\mathbf{d}_{k}$ is a vector of pairwise inverse distances between units, based on their spatial encoding, and $\mathrm{corr}$ is Pearson’s $r$. This means that nearby units (i.e. high inverse distance) should have highly correlated activations on the same inputs, and that distant units should be less correlated; this gives us a notion of spatial smoothness. We scale by a factor of $0.5$ to ensure that $\mathrm{SL}\in[0,1]$.
We compute this spatial loss for every attention and MLP layer in a Transformer, prior to normalization and addition into the residual stream. Rather than computing the spatial loss for the entire layer, as in Margalit et al. (2024) we approximate the loss using small neighborhoods, ensuring that the model optimizes for local, rather than global ‘long-distance’ constraints. [^3] This could be controlled for using many alternative methods, e.g. a Gaussian smoothing kernel. However, we choose to use neighborhood-level approximations for simplicity. For each batch of inputs, the model is then optimized subject to the loss criterion
@@
In the below experiments, we utilize an adapted GPT-2-small style architecture (Radford et al., 2019). We use hidden dimension 784 such that we can evenly embed units in a $28\times 28$ grid. The model has 12 Transformer blocks, each with 16 attention heads and a GELU activation function. We train our models on a randomly sampled 10B-token subset of the FineWeb-Edu dataset. The task loss is cross-entropy on next-word prediction. We use batch size 48 and block size 1024. For spatial loss, we set $\alpha_{k}=2.5$ across all layers [^4] We chose this value of $\alpha$ after extensive hyperparameter search. In particular, lower values of $\alpha$ do not adequately encourage the development of topography, while greater values impede task performance and the development of meaningful representations. and operationalize the inverse distance vector $\mathbf{d}$ with the $\ell^{\infty}$ norm. For each batch, we average the spatial loss across $5$ randomly selected neighborhoods, each of $\ell^{\infty}$ radius $5$. This allows us to compute loss more efficiently without significant performance drops.
@@
In all below analyses, we compare TopoLM to BinHuraib et al. (2024)’s pre-trained Topoformer-BERT, a BERT-style model (Devlin et al., 2019) with local connectivity trained on the BookCorpus dataset (Zhu et al., 2015). [^6] Topoformer-BERT has $16$ layers, each with a single attention head, with topographic constraints applied both at the output of each attention layer and after the key matrix product; we therefore compare TopoLM to outputs at these levels. We refer to BinHuraib et al. (2024) for more details. Note critically that Topoformer-BERT is a baseline, but not a control—it is trained on a much smaller corpus, has only one attention head per layer, and is bidirectional.
Figure 2: Brain-like response profiles across the core language system. (a) Applying a functional localizer (Fedorenko et al., 2010) we isolate the core language system of TopoLM, and find clear brain-like spatial organization (for brevity, we only show Transformer blocks 5-12 here). Response profile across individual language-selective clusters in TopoLM are similar to one another, consistent with (b) the language system in human cortex (Fedorenko et al., 2024). (c) Across the entire core language system, TopoLM (blue) mostly matches the neural data (green), but not exactly; however, the non-topographic baseline model (orange) fails to capture neural patterns as well.
@@ 4 Spatio-Functional Organization of the Core Language System
Language processing in the brain engages a set of left-lateralized frontal and temporal brain regions. These areas are typically referred to as the ’core language system’ (Fedorenko et al., 2010) and respond selectively to linguistic input in contrast to non-linguistic stimuli (see Fedorenko et al., 2024, for an overview). Due to anatomical differences between individuals, the language system is defined via a functional localizer that contrasts syntactically and semantically valid sentences against a perceptually matched control, such as strings of nonwords (Fedorenko et al., 2011).
@@
Prior work on neural alignment in language models typically compares neural responses across the entire core language system of the brain to model activations. The topographic organization of our model enables us to test for the emergence of a brain-like spatially organized core language system in silico. A successful spatio-functional alignment between brain and model would mean that (1) distinct language-selective clusters emerge in the model, (2) these clusters all have consistent response profiles similar to consistent response profiles across sub-regions of the ’core language system’ in humans 777”The language areas all show a similar response profile (despite slight apparent differences, no region by condition interactions come out as reliable, even in well-powered studies.” Fedorenko et al. (2024), , and (3) the response profiles match the activity profiles in the brain (sentences >>> {unconnected words, jabberwocky} >>> nonwords; AlKhamissi et al., 2024).
@@ Methods.
To isolate the core language system in TopoLM, we use the same localization stimuli as Fedorenko et al. (2010), which consists of a set of 160 sentences and 160 strings of non-words, all 12 words each. After passing these through the model, at each attention and MLP layer we run a $t$-test across the activations of all layer units. We then define the core language system as all units that are significantly language-selective ($p<0.05$ after correction for multiple comparison across all layers using the false-discovery-rate (FDR) (Benjamini & Hochberg, 1995)).
We then define language-selective clusters using an evolutionary clustering algorithm applied to each contrast map. In each layer, we begin with the most selective unit by $t$-value, and then repeatedly add the most selective neighboring unit to the cluster, until we hit a pre-determined $p$-value threshold associated with the unit’s selectivity (here, $p(\text{FDR})<0.05$). We repeat this process, searching for new clusters until all units in the layer are exhausted. We discard clusters with fewer than $10$ units.
@@ Results.
TopoLM exhibits clear brain-like spatial organization of the language network, such that (1) multiple language-selective clusters emerge across the topographic tissue (Figure 2A and Figure 8), (2) across most clusters, the response profiles are consistent with one another (Figure 9), and (3) response profiles mostly match the ones in the brain. The response profiles are not a perfect match to the brain data—while sentences have higher activations than Jabberwocky and nonword stimuli, they do not have higher activation than unconnected words as in brain data. However, looking across the entire language selective network, the response profile of the non-topographic baseline model similarly fails to capture the neural response profile (Figure 2C), suggesting a general shortcoming of the base transformer model, rather than a weakness of topography. We similarly find evidence for language-selective clustering in Topoformer-BERT (see Figure 13).
5 Spatio-Functional Organization Of Semantic Clusters
Beyond selectivity for language in general, experimental evidence supports the existence of cortical noun- and verb-selective clusters across human subjects during processing of verbal and nominal stimuli in auditory (Elli et al., 2019; Hauptman et al., 2024), visual (Moseley & Pulvermüller, 2014), and speech production (Shapiro et al., 2006) tasks. Here, we compare spatial activation patterns predicted by TopoLM to two groups of fMRI studies: Elli et al. (2019) / Hauptman et al. (2024), who use the same set of stimuli and the same experimental setup (Figure 3 and Appendix Figures 7 and 10), and Moseley & Pulvermüller (2014) (Figure 4). We evaluate TopoLM predictions quantitatively using Hauptman et al. (2024)’s fMRI data [^8] Available on OPENICPSR at https://doi.org/10.3886/E198163V3. and perform qualitative evaluations where no neural data was available (Elli et al., 2019; Moseley & Pulvermüller, 2014).
Figure 3: Brain-like verb- and noun-selective clusters in TopoLM. (a) fMRI data from Hauptman et al. (2024) points to verb- and noun-selective regions in the human cortex with strong clustering (Moran’s $I=0.96$). (b) Quantification of clustering. Relative to high clustering in the brain (green dashed line), the non-topographic baseline shows limited clustering (orange). The topographic model shows moderate clustering at the unit level (light blue) and strong clustering when simulating fMRI sampling (dark blue). On stimuli from Moseley & Pulvermüller (2014) (fMRI data not available) we find qualitatively similar results. (c) Exemplary model maps (last MLP layer) showing the verb-/noun contrast (red-blue) in response to stimuli from Hauptman et al. (2024). The non-topographic baseline shows no clustering while the topographic model develops verb- and noun-selective clusters.